OpenStudy (curry):

((x-1)/(x^2-2x)) + ((1)/(2x-4)) ((x-1)/(x^2-2x)) + ((1)/(2x-4)) @Mathematics

OpenStudy (anonymous):

\(\frac{x-1}{x(x-2)}+\frac{1}{2(x-2)}=\frac{2(x-1)}{2x(x-2)}+\frac{x}{2x(x-2)}=\frac{3x-2}{2x(x-2)}.\)

hero (hero):

Curry you don't really need help with this